OpenStudy (anonymous):

Algebra 2 help? Adding and subtracting radical expressions... So lost

OpenStudy (anonymous):

i got a 100% on that

OpenStudy (anonymous):

Wanna help me out?? :)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

\[7\sqrt{13x}-15\sqrt{2x}+2\sqrt{13x}-4\sqrt{2x}\]

OpenStudy (anonymous):

like asking what is \[7y-15z+2y-4z\] exactly

OpenStudy (anonymous):

Just combine like terms

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

for example \(7\sqrt{13x}+2\sqrt{13x}=9\sqrt{13x}\)

OpenStudy (anonymous):

and \[-15\sqrt{2x}-4\sqrt{2x}=-19\sqrt{2x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

put them together and get \[9\sqrt{13x}-19\sqrt{2x}\] there is nothing more than can be done

OpenStudy (anonymous):

since \(\sqrt{13x}\)and \(\sqrt{2x}\) are not like terms, you cannot combine them in one term

OpenStudy (anonymous):

So that is all? Wow, that's much easier than I thought! Thank you so much! ;D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

like most math, it is easy when you know what to do